Use the substitution method to solve the system of equations. Enter your answer as an ordered pair.
3x - y = 7
2x - 2y = 2

Found 2 solutions by checkley77, rfer:
Answer by checkley77(12844)

3x - y = 7 multiply by -2 & add
2x - 2y = 2
-6x+2y=-14
------------------
-4x=-12
x=-12/-4
x=3 ans.
3*3-y=7
9-y=7
-y=7-9\-y=-2
y=2 ans.
Proof:
2*3-2*2=2
6-4=2
2=2

Answer by rfer(16322)

3x-y=7, y=3x-7
2x-2y=2
2x-2(3x-7)=2
2x-6x+14=2
-4x=-12
x=3
----------------
3(3)-y=7
-y=-2
y=2
---------------
2(3)-2(2)=2